OpenStudy (anonymous):

I need Help fast!!!!! Fan and Medal!!!! Is it possible to have a rectangular playground area made from 70 yards of fencing with an area of 324 square yards? Explain with calculations.

OpenStudy (anonymous):

Not if it's rectangular. The max area is a square --> sides = 17.5 yards Area = 17.5^2 = 306.25 sq yds If it's circular: C = 2pi*r = 70 r = 35/pi yds Area = pi*r^2 = pi*1225/pi^2 = 1225/pi sq yds Area =~ 389.93 sq yds

OpenStudy (anonymous):

Is it possible to have a playground area made from 70 yards of fencing with an area of 324 square yards? Explain with calculations